1. Given f(x,y) = z as z = 2 +y find: (a) the partial derivative f(x,y). (b) the partial derivative fy(2,y).
Answer #1

Let 2= I zaty a fixe (2x+422 filme) = -2 (29448 6 fy (x,y)= a ( Fenw) -1 9 (20+4J Canty 2 y - 1 1 cox tyg2 - fylx,y) = ale -